A car (mass with a load of 1.8 t) drives at a steady speed up a slope, whose slope angle is 0.12.
The various movement resistances together account for 8.0% of the car's weight. Calculate the car's speed when moving
Maintenance requires a power of 25 kW. (Answer 26 km / h)
Answer #1

mass m= 1.8 tonene = 1.8 tonne m = 1.8x1000 kg = 1800 kg 0 = 0.12° = angle of Inclination. Resistance R=8% of weight R= . (19from free body diagram of car: W sino < 10--2--+ N= noimal force. f = applied external force W= weight f = friction . appliedF = (18000) sin (0.12 x 150) + 1440 2154.82 + 1440 3595 N. lelocity v= p. 25000 - 3595 V= 6.95 m/s V= 6.95 x 15 kmph V=25 kmp